From mar@roe.ac.uk Thu Aug 14 08:46:25 2014 Date: Thu, 14 Aug 2014 08:37:23 +0100 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Dave Morris , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Alastair Edge , Andrew Lawrence , Bob Mann , Dante Minniti , James Dunlop , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Maria-Rosa Cioni , matt.jarvis@astro.ox.ac.uk, Mike Irwin , Nigel Metcalfe , Norman Gray , Phil Lucas , Richard McMahon at IOA , Simon Dye , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives progress meeting minutes, 8 August 2014 WFAU VDFS Science Archives progress meeting minutes, 8 August 2014 Present: RSC, ETWS, MAR, ACD, MSH Apologies: RPB, NJC, RGM, NCH, KTN, AL, STV DoNM: 10am Friday 22 August 2014 in the Villas meeting room, MAR in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: MAR to document issue with ATLAS unpaired sources and give examples of recovering matches. Discharged, published on OSA under Known Issues. ACTION NJC: update ESO VISTA releases on the deadlines web page. Discharged. ACTION: ACD ensure that the Twiki deadline page is updated for GES iDR3. Discharged. Actions partly discharged but continuing: ---------------------------------------- None. Actions carried forward from 20/6/2014 meeting: ---------------------------------------------- ACTION: RPB to put all prices of additional items associated with the backup policy on the Twiki pages. Continues. RPB had obtained the prices but not yet put them on the Twiki page. In summary the costs would amount to a one-off expense of 7000 UKP and a recurrent monthly expense of 130 UKP. ACTION: RGM to investigate the size and content (how does it differ from DR9?) of DR10 to see if it's worth us hosting and whether it is necessary to keep all of the other releases. Continues. RGM had started a wiki page for WFAU SDSS strategy http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. ACTION: ACD to review VVV brainstorming Twiki page and decide who to chase about progressing their assigned actions. Continues: ACTION: RPB to write a script to backup the SVN repository. Continues ACTION: MSH to produce a costed list of the additional kit necessary to permit wholesale reprocessing of UKIDSS/VISTA. Continues. ACTION: ACD, RGM: review finances. Continues. ACTION: RPB to make the necessary Spitzer SMC/VMC neighbour table entries. Continues (RPB not present). Specific points and new actions: ------------------------------- Project management: Nothing to report. WFCAM, VISTA and VST updates: Nothing to report. Comments and issues arising from CASU minutes: No minutes had been received since the previous meeting. Review of Deadlines: ACD took the meeting through the relevant upcoming deadlines. Most seemed to be on track, see notes below. 15/8 - VMC P92 release was still running and would likely be a few days late. 29/8 - GPS release into DR10, still the goal but had just re-transferred and re-ingested some data so not yet running. 25/9 - start CU19 for VVVDR3 - this would need to be pushed back as objID updates were still being made and additional QC needed to be applied. A more realistic target is to kick off CU19 in December. Networking: Up and running: WSA/VSA/OSA/GES Operations: ETWS reported that: - GPS re-transfers had been done and all GPS data for the DR10 release were now ingested. - OSA (v1.1) CU3 of P92 was almost complete. - ObjID updates were still running on VVV. - VMC was currently ingesting the results of the deep stacking NJC reported that he had downloaded the latest VIDEO files from the consortium. ACD noted that we were awaiting revisions to WG15 data and that these would be inserted into GESiDR2. Hardware and Systems: MSH reported that:# - the new database node was installed and ready for use. - the 10GB ethernet switch had arrived. The network would be set-up over the coming weeks Software: NJC had been working on the stacking code and the synoptic set-up for VVV. ETWS had made some updates to CU4. Survey Data Release: We were still iterating with ESO over the latest VIKING submission. Non-survey Data Release: No non-survey releases had been made since the last meeting. ACTION: MAR to have a look at the status of non-survey releases for registered projects. Astrogrid deployment, VO & Data Analysis services: STV had set running some intensive tests of the new UI infrastructure. Miscellaneous: ACD reminded the meeting of the ROE open days, Sept 27 & 28. -- Scanned by iCritical.